# Strategic resource allocation in memory encoding: An efficiency principle shaping language processing

### v1

@@ 6.6 Surprisal effect as efficient coding
Here is one way to think about the basic surprisal effect from an information-theoretic perspective. More surprising linguistic units correspond to longer code length. For example, consider a low-surprisal word encoded by a sequence of 3 bits ‘110’, compared to a high-surprisal word is encoded by 11 bits ‘11100001101’. A simple mechanical interpretation of the basic surprisal effect, therefore, is that encoding longer sequence of code in memory requires more time and effort. This leads to the widely observed linear relationship between surprisal and behavioral measures, such as reading time (e.g., Smith & Levy, 2013; Hoover et al., 2023; Shain et al., 2024; Wilcox et al., 2023; Xu et al., 2023).
@@
However, choosing a code length based on predictability is not the only possible strategy to encode a linguistic unit into memory. An alternative approach is to assign each unit a code of equal length regardless of statistical regularities, a scheme analogous to the uniform distribution of resource allocation discussed in the Introduction. For example, the ASCII (American Standard Code for Information Interchange) system uses such a coding strategy, where every character is represented as a 7-bit sequence. Under this hypothetical scenario, processing effort would be insensitive to statistical regularities and uniformly distributed across all linguistic units.
@@
As demonstrated in previous work on information theory, a coding scheme that treats every input as equally likely is inherently inefficient. When the statistical structure of the input is known, a coding scheme that assigns shorter code lengths to more predictable inputs can reduce the average code length, thus increasing efficiency (Shannon, 1948). The observed linear relationship between surprisal and processing effort suggests that the brain may adopt such an efficient coding strategy. Specifically, the brain appears to assign code lengths to input units in proportion to their likelihood of occurrence based on long-term statistical regularities. This strategy is supported by the evidence that the brain is very good at inferring and approximating the statistical structure of the external environment (Saffran et al., 1996), and there is good reason to believe that the brain uses these inferred statistical structure to encode information more efficiently.
@@
The connection between the surprisal effect and efficient coding suggests that the basic surprisal effect by itself may be the result of evolutionary adaptation to efficiency pressures. Under these pressures, the cognitive system may have evolved to adopt a memory encoding strategy that optimizes an objective function within the constraints of limited resources. The exact nature of this objective function remains an open question: it could be the minimization of distortion cost, or it could involve something else. The key point, however, is that the relationship between surprisal and strategic resource allocation can be understood at different timescales. On the one hand, at the level of processing individual sentences, strategic resource allocation may arise as a by-product of the effort required to process surprising information. On the other hand, over a longer timescale, the tendency to invest more effort in encoding less predictable information may reflect an evolved strategy that is adapted to an efficiency problem.
@@ 6.7 Limitations
One major limitation of our analysis is that it heavily depends on the quality of the surprisal measures generated from LLMs. In the current study, we conducted our analysis using the surprisal from two language models, namely GPT-3 and mGPT. As presented above, the two models do yield consistent pattern, alleviating the concern that our main result may be the artifact of any model-specific behavior. However, it still does not entirely rule out the issue with the accuracy of LLM surprisals, especially for low-resource languages that are under-represented in the training data of the models we used. This limitation may compromise the extensibility of our analysis to under-studied languages, which are of particular interest from a typological perspective, and is particularly relevant for the unreliable effect we observed for some non-Indo-European languages.
@@ 7 Conclusion
The current study investigates strategic resource allocation as an efficiency principle for memory encoding in sentence processing, which holds that working memory resources are strategically and dynamically allocated to prioritize novel and unexpected information. Empirically, this principle predicts that the memory representation of more surprising linguistic units is more robust against interference and decay. We examined this prediction through naturalistic corpus data in the context of dependency locality from both the comprehension and the production side. In production, through the analysis of UD corpora in 10 languages, we indeed found that more surprising antecedents can tolerate longer dependency length, but the effect mostly exists within Indo-European and head-initial languages. This cross-linguistic variability therefore calls for a closer look into the interaction between strategic resource allocation as a universal efficiency principle and language-specific phrase structure in future work. In comprehension, through two English reading time corpora, we observed a similar reduced locality effect on retrieval difficulty for more surprising antecedents. Moreover, in all analyses, we found that the effect is more reliable for dependencies of subject relations than object relations. Taken together, there is converging evidence from naturalistic corpus data supporting that unpredictable antecedents are encoded with enhanced representation to be more resistant against memory decay and interference, a pattern that is predicted by strategic resource allocation. For linguistic typology, our findings contribute to the functionalist approach, highlighting the need to examine more sophisticated functional constraints in terms of their explanatory power for the structure of human language.

### v2

@@ 7.8 Surprisal effect as efficient coding: An adaptionist view
Here is one way to think about the basic surprisal effect from an information-theoretic perspective. More surprising linguistic units correspond to longer code length. For example, consider a low-surprisal word encoded by a sequence of 3 bits 110, compared to a high-surprisal word encoded by 11 bits 11100001101. An (over-)simplified mechanical interpretation of the basic surprisal effect, therefore, is that encoding longer sequence of code in memory requires more time and effort. This leads to the widely observed linear relationship between surprisal and behavioral measures, such as reading time (e.g., Smith & Levy, 2013; Hoover et al., 2023; Shain et al., 2024; Wilcox et al., 2023; Xu et al., 2023).
@@
However, choosing a code length based on predictability is not the only possible strategy to encode a linguistic unit. An alternative approach is to assign each unit a code of equal length regardless of statistical regularities, a scheme analogous to the uniform distribution of resource allocation discussed in Introduction. For example, the ASCII (American Standard Code for Information Interchange) system uses such a coding strategy, where every character is represented as a 7-bit sequence. Under this hypothetical scenario, processing effort would be insensitive to statistical regularities and uniformly distributed across all linguistic units.
@@
As demonstrated in previous work on information theory, a coding scheme that treats every input as equally likely is inherently inefficient. When the statistical structure of the input is known, a coding scheme that assigns shorter code lengths to more predictable inputs can reduce the average code length, thus increasing efficiency (Shannon, 1948). The observed linear relationship between surprisal and processing effort suggests that the brain may adopt such an efficient coding strategy. Specifically, the brain appears to assign code lengths to input units in proportion to their likelihood of occurrence based on long-term statistical regularities. This strategy is supported by the evidence that the brain is very good at inferring and approximating the statistical structure of the external environment (Saffran et al., 1996), and there is good reason to believe that the brain uses these inferred statistical structures to encode information more efficiently.
@@
The connection between the surprisal effect and efficient coding suggests that the basic surprisal effect itself may be construed in evolutionary terms. Under certain efficiency pressures, the cognitive system may have evolved to adopt a memory encoding strategy that optimizes an objective function within the constraint of limited resources. The exact nature of this objective function remains an open question: it could be the minimization of distortion cost, or it could involve something else. The key point, however, is that the relationship between surprisal and strategic resource allocation can be understood at different timescales. On the one hand, at the level of processing individual sentences, strategic resource allocation may arise as a by-product of the effort required to process surprising information. On the other hand, over a longer timescale, the tendency to invest more effort into encoding less predictable information may reflect an evolved strategy that is adapted to an efficiency problem.
@@ 7.9 Limitations
One major limitation of our analysis is that the results heavily depend on the quality of surprisal measures generated from LLMs. In the current study, we conducted our analysis using the surprisal from two language models, namely GPT-3 and mGPT. As presented above, the two models do yield consistent pattern, alleviating the concern that our main result may be the artifact of any model-specific behavior. However, it still does not entirely rule out the issue with the accuracy of LLM surprisals, especially for low-resource languages that are under-represented in the training data of the models we used. This limitation may compromise the extensibility of our analysis to under-studied languages, which are of particular interest from a typological perspective, and is particularly relevant for the unreliable effect we observed for some non-Indo-European languages.
@@ 8 Conclusion
The current study proposes Strategic Resource Allocation (SRA) as an efficiency principle for memory encoding in sentence processing, which holds that working memory resources are strategically and dynamically allocated to prioritize novel and unexpected information. Theoretically, we argue that SRA is an efficient solution to the computational problem faced by working memory, that is, to maximize the retrieval accuracy of past inputs under the constraint of limited memory resources. Empirically, this principle predicts that the memory representation of more surprising linguistic units is more robust against interference and decay. We examined this prediction through naturalistic corpus data in the context of dependency locality from both the comprehension and the production side. In production, through the analysis of UD corpora in 10 languages, we indeed found that more surprising antecedents can tolerate longer dependency length, but the effect mostly exists within Indo-European and head-initial languages. This cross-linguistic variability therefore calls for a closer look into how SRA as a domain-general memory efficiency principle interacts with the language-specific phrase structure. In comprehension, through two English reading time corpora, we observed a similar reduced locality effect on retrieval difficulty for more surprising antecedents. Moreover, we found that the effect is more reliable for dependencies of subject relations than object relations. Taken together, there is converging evidence from naturalistic corpus data supporting that unpredictable antecedents are encoded with enhanced representation to be more resistant against memory decay and interference, a pattern that is predicted by our SRA.
Appendix A Mathematical derivation of strategic resource allocation
In this section, we demonstrate the mathematical derivation of strategic resource allocation. First, we will characterize the memory retrieval process as Bayesian inference for the encoded linguistic input (e.g., words). For the purpose of this derivation, we will assume that the linguistic prediction and the underlying memory representation follow Gaussian distributions; these may be interpreted as distributions over values of features. Needless to say, this is a highly simplified view of mental lexicon, and is not necessarily the reality for memory encoding, especially given that word inputs are discrete units rather than continuous variables. However, Gaussian distribution has some desirable mathematical properties with analytical solutions to help us validate the intuition behind our proposal.
A.1 Memory retrieval via Bayesian inference
Suppose one is trying to encode an input word $w$ in noisy memory. We model noise by assuming that the input representation $w$ is corrupted by Gaussian noise with an adjustable precision $\tau_{w}$, yielding a noisy memory representation $r$:
$r\sim w+\mathcal{N}\left(0,\tau_{w}^{-1}\right).$ (A.1)
Retrieval from memory is then performed by forming a reconstructed representation $\hat{w}=\mathop{\mathbb{E}}_{p(w\mid r)}\left[w\right]$ as the posterior mean on input representations $w$ given noisy memory representations $r$ and a prior distribution on inputs $p_{0}$, with posterior distribution
$p(\hat{w}\mid r)\propto p_{M}(r\mid w)p_{0}(w).$ (A.2)
Now for ease of analysis, we set the prior distribution on input representations $w$ to be a Gaussian distribution parameterized with mean $w_{0}$ and the precision $\tau_{0}$:
$p_{0}(w)=\mathcal{N}\left(w\mid w_{0},\tau_{0}^{-1}\right).$ (A.3)
A useful property of Gaussian distribution is that it forms a conjugate prior, so the posterior distribution on inputs words $w$ given memory representations $r$ is also Gaussian distribution. We assume that, in memory retrieval, the decoder  [^18] This assumes that the decoder has access to the encoding precision $\tau_{w}$ for the input.
$\hat{w}\mid r\sim\mathcal{N}\left(\mu_{\mathrm{post}},\tau_{\mathrm{post}}^{-1}\right),$ (A.4)
where the posterior mean and precision are
$\mu_{\mathrm{post}}=\left(1-\alpha_{w}\right)w_{0}+\alpha_{w}r,\text{ }\tau_{\mathrm{post}}=\tau_{0}+\tau_{w},$ (A.5)
where $\alpha_{w}=\frac{\tau_{w}}{\tau_{0}+\tau_{w}}$. Then marginalizing out the memory representations, the distribution on reconstructed words is
$\hat{w}\mid w\sim\mathcal{N}\left(\alpha_{w}w+\left(1-\alpha_{w}\right)w_{0},\frac{\alpha_{w}}{\tau_{0}+\tau_{w}}\right).$ (A.6)
We see that the retrieved word $\hat{w}$ is pulled towards the prior mode $w_{0}$ with a weight that depends on the encoding precision $\tau_{w}$. As the encoding precision $\tau_{w}$ increases, this attraction to the prior is reduced.
Expected retrieval error under the memory model.
We define the expected retrieval error $\varepsilon(w)$ for input $w$ as the mean squared error between input $w$ and reconstruction $\hat{w}$:
$\varepsilon(w)=\mathop{\mathbb{E}}\left[\left(\hat{w}-w\right)^{2}\mid w\right].$ (A.7)
We can express this mean squared error in terms of the bias–variance decomposition as
$\varepsilon(w)=\mathrm{Var}[\hat{w}\mid w]+\left(\mathop{\mathbb{E}}\left[\hat{w}\mid w\right]-w\right)^{2}.$ (A.8)
Dropping in the mean and variance from Eq. A.6, we can express the bias and variance as
$\mathrm{Var}\left[\hat{w}\mid w\right]$ $=\frac{\alpha_{w}}{\tau_{0}+\tau_{w}}$ (A.9)
$\mathop{\mathbb{E}}\left[\hat{w}\mid w\right]-w$ $=\alpha_{w}w+\left(1-\alpha_{w}\right)w_{0}-w$ (A.10)
$=\left(1-\alpha_{w}\right)\left(w_{0}-w\right).$ (A.11)
This gives us a convenient expression for the expected retrieval error,
$\varepsilon(w)$ $=\frac{\tau_{w}}{\left(\tau_{w}+\tau_{0}\right)^{2}}+\left(\frac{\tau_{0}}{\tau_{0}+\tau_{w}}\left(w_{0}-w\right)\right)^{2}$ (A.12)
$=\frac{\tau_{w}+\tau_{0}^{2}\left(w-w_{0}\right)^{2}}{\left(\tau_{0}+\tau_{w}\right)^{2}}.$ (A.13)
Furthermore, we will wish to express the expected retrieval error in terms of the surprisal $h_{w}=-\ln p_{0}(w)$ of input $w$ and the encoding precision $\tau_{w}$. From the assumed Gaussian form of the prior over input words, we have
$p_{0}(w)$ $=\sqrt{\frac{\tau_{0}}{2\pi}}\exp\left(-\frac{\tau_{0}(w-w_{0})^{2}}{2}\right)$ (A.14)
$h_{w}$ $=\frac{\tau_{0}(w-w_{0})^{2}}{2}-\ln\sqrt{\frac{\tau_{0}}{2\pi}}.$ (A.15)
Extracting $\tau_{0}(w-w_{0})^{2}$ to the left-hand side yields:
$\tau_{0}(w-w_{0})^{2}=2h_{w}+\ln\frac{\tau_{0}}{2\pi}.$ (A.16)
Substituting this back into the expression for expected retrieval error, we get
$\varepsilon(h_{w},\tau_{w})=\frac{\tau_{w}+2\tau_{0}h_{w}+\tau_{0}\ln\frac{\tau_{0}}{2\pi}}{\left(\tau_{0}+\tau_{w}\right)^{2}}.$ (A.17)
Below, we will consider how to choose encoding precisions in order to minimize the expected retrieval error under constraints.
A.2 Minimizing expected error under memory constraint
In this section, we aim to show that strategic resource allocation arises as a solution to the problem of minimizing the expected retrieval error on average:
  • Strategic Resource Allocation
    Given two linguistic inputs, the minimization of their total expected error bounded by certain memory constraint requires that the more surprising input be encoded with higher precision.
Consider two input words $w_{1}$ and $w_{2}$ to be encoded and retrieved, whose surprisals are $h_{w_{1}}=-\ln p_{0}(w_{1})$ and $h_{w_{2}}=-\ln p_{0}(w_{2})$ respectively. We assume that the encoding precision $\tau_{w}$ for each word is proportional to the memory resources allocated, and we assume that there is a constraint on total memory resources $c$ allocated for both words, which is to be distributed between $w_{1}$ and $w_{2}$. That is, we posit a constraint on the the sum of encoding precisions $\tau_{w_{1}}+\tau_{w_{2}}=c$.
We will show that this optimization problem bounded by memory constraint leads to strategic resource allocation in memory encoding, as stated in the Proposition below:
Proposition.
To minimize the total expected retrieval error for two linguistic inputs $\varepsilon_{w_{1}}+\varepsilon_{w_{2}}$ subject to a resource constraint $\tau_{w_{1}}+\tau_{w_{2}}=c$, the input that is more surprising under the prior distribution must be encoded with higher precision. Specifically, if
$h_{w_{1}}>h_{w_{2}},$ (A.18)
then the optimal encoding satisfies
$\tau_{w_{1}}>\tau_{w_{2}}.$ (A.19)
Proof.
First, in order to minimize expected error $\varepsilon$, we take the derivative of $\varepsilon$ with respect to encoding precision $\tau_{w}$ for each input [^19] For simplicity, we maintained $(w-w_{0})^{2}$ for now in Eq. A.20 instead of having it transformed to the form that contains the surprisal $h_{w}$ of the input.:
$\frac{\partial\epsilon(\tau_{w},h_{w})}{\partial\tau_{w}}=\frac{\tau_{0}-\tau_{w}-2\tau_{0}^{2}\left(w-w_{0}\right)^{2}}{\left(\tau_{0}+\tau_{w}\right)^{3}},$ (A.20)
which reveals three possible situations with respect to the monotonicity of $\varepsilon$ for both inputs:
  1. 1.
    The expected error monotonically decreases with increasing $\tau_{w}$ within its meaningful domain (i.e., $\tau_{w}>0$) for both inputs $w_{1}$ and $w_{2}$. As shown below, this is the situation for most cases where the input $w$ is not too close to the prior prediction $w_{0}$ and the prior precision $\tau_{0}$ is not too unreliable.
  2. 2.
    The expected error is a non-monotonic function of $\tau_{w}$ for both inputs $w_{1}$ and $w_{2}$.
  3. 3.
    The expected error monotonically decreases with $\tau_{w}$ within in meaningful domain for one input but is non-monotonic for the other.
In this proof, we will show that the proposition above holds in all these three situations.
Situation 1.
In this first situation, the expected error $\varepsilon$ monotonically decreases for both inputs, which means that
$\frac{\partial\epsilon(\tau_{w},h_{w})}{\partial\tau_{w}}\leq 0.$ (A.21)
Thus, Situation 1 holds when
$\tau_{0}-\tau_{w}-2\tau_{0}^{2}\left(w-w_{0}\right)^{2}\leq 0$ (A.22)
$\tau_{w}\geq\tau_{0}-2\tau_{0}^{2}\left(w-w_{0}\right)^{2}.$ (A.23)
Since $\tau_{w}>0$, the inequality in Eq. A.23 always hold within the meaningful domain of $\tau_{w}$ if
$\tau_{0}-2\tau_{0}^{2}\left(w-w_{0}\right)^{2}\leq 0$ (A.24)
$\tau_{0}\geq\frac{1}{2\left(w-w_{0}\right)^{2}}.$ (A.25)
Intuitively, this means that the expected error $\varepsilon$ monotonically decreases if the input $w$ is not too close to the prior prediction $w_{0}$ and the prior precision $\tau_{0}$ is not too unreliable.
Now that $\varepsilon$ monotonically decreases within the meaningful domain of $\tau_{w}$ for both inputs $w_{1}$ and $w_{2}$ as defined in this first situation, using Eq. A.17, we evaluate the derivative with respect to its relationship with input surprisal
$\begin{split}f(\tau_{w},h_{w})&=\frac{\partial\epsilon(\tau_{w},h_{w})}{\partial\tau_{w}}\\ &=\frac{\tau_{0}-\tau_{w}-2\tau_{0}\left(2h(w)+\ln\left(\frac{\tau_{0}}{2\pi}\right)\right)}{(\tau_{0}+\tau_{w})^{3}}\\ &<0\text{ (by the definition of Situation 1)},\end{split}$ (A.26)
which shows that $f(\tau_{w},h_{w})$ monotonically decreases in $h_{w}$.
When $h_{w_{1}}=h_{w_{2}}$, $f(\tau_{w_{1}},h_{w_{1}})$ and $f(\tau_{w_{2}},h_{w_{2}})$ have the same quantitative form. As a result, the memory resources will be uniformly distributed across $w_{1}$ and $w_{2}$ with no impetus to redistribute more resources to any one of them:
$\tau_{w_{1}}=\tau_{w_{2}}=\frac{c}{2}.$ (A.27)
However, when $h_{w_{1}}>h_{w_{2}}$, since $f(\tau_{w},h_{w})$ monotonically decreases in $h_{w}$, we have $f(\tau_{w_{1}},h_{w_{1}})$ < $f(\tau_{w_{2}},h_{w_{2}})$ at any given value of $\tau_{w}$. As a result, compared to the uniform distribution in Eq. A.27, there is reason to redistribute more resources to to the high surprisal $w_{1}$, since the decrease of error on $w_{1}$ will be higher than the increase of error on $w_{2}$, yielding a lower total error across the two inputs. Therefore, in Situation 1, if $h_{w_{1}}>h_{w_{2}}$, the optimal encoding strategy should satisfy $\tau_{w_{1}}>\tau_{w_{2}}$.
Figure A1: Situation 1 expected retrieval error $\epsilon$ for two input words as a function of encoding precision $\tau_{w}$ and their surprisal.
Situation 2.
As shown above, in order for $\varepsilon$ to be a non-monotonic function of $\tau_{w}$:
$\tau_{0}<\frac{1}{2\left(w-w_{0}\right)^{2}},$ (A.28)
which corresponds to borderline cases where the input $w$ is too close to the prior prediction $w_{0}$ or the prior precision $\tau_{0}$ is too unreliable.
In this second situation, with increasing $\tau_{w}$, $\varepsilon$ first increases and then decreases, as illustrated in Figure A2. And the relationship between high surprisal and low surprisal inputs has three phases. We will show that the proposition still holds for all the three phases in Situation 2, such that the optimal encoding satisfies $\tau_{w_{1}}>\tau_{w_{2}}$ if $h_{w_{1}}>h_{w_{2}}$.
Figure A2: Situation 2 expected retrieval error $\epsilon$ for two input words as a function of encoding precision $\tau_{w}$ and their surprisal.
In Phase III, the situation is basically the same as the Situation 1 discussed above, where $\varepsilon$ decreases with increasing $\tau_{w}$ for both inputs. Therefore, according to the proof in Situation 1, the optimal encoding satisfies $\tau_{w_{1}}>\tau_{w_{2}}$ if $h_{w_{1}}>h_{w_{2}}$.
In Phase II, $\varepsilon$ decreases in $\tau_{w}$ for one input and increases for the other. For each input, the turning point where the monotonicity is flipped is at
$\tau_{w}=\tau_{0}-2\tau_{0}\left(2h(w)+\ln\left(\frac{\tau_{0}}{2\pi}\right)\right),$ (A.29)
where the derivative of $\varepsilon$ in Eq. A.26 is 0. Importantly, if $h_{w_{1}}>h_{w_{2}}$, then the turning points $\tau_{w_{1}}<\tau_{w_{2}}$. Therefore, the turning point for the high surprisal input $w_{1}$ is to the left of the one for the low surprisal input $w_{2}$. That means, in Phase II, it can only be the case that the expected error $\varepsilon$ decreases in $\tau_{w}$ for the high surprisal $w_{1}$, but increases for the low surprisal $w_{2}$. As a result, in order to minimize $\varepsilon$, more memory resources should be allocated to encode $w_{1}$ than $w_{2}$, leading to $\tau_{w_{1}}>\tau_{w_{2}}$.
In Phase I, $\varepsilon$ increases in $\tau_{w}$ for both inputs $w_{1}$ and $w_{2}$. Recall that the derivative of $\varepsilon$ (repeated below in Eq. A.30) decreases as the input surprisal $h_{w}$ increases.
$\begin{split}f(\tau_{w},h_{w})&=\frac{\partial\epsilon(\tau_{w},h_{w})}{\partial\tau_{w}}\\ &=\frac{\tau_{0}-\tau_{w}}{(\tau_{0}+\tau_{w})^{3}}-\frac{2\tau_{0}\left(2h(w)+\ln\left(\frac{\tau_{0}}{2\pi}\right)\right)}{(\tau_{0}+\tau_{w})^{3}}\\ &>0.\end{split}$ (A.30)
Therefore, in Phase I, the expected error $\varepsilon$ increases more slowly for high surprisal input. As a result, if $h_{w_{1}}>h_{w_{2}}$, then for a fixed amount of memory resources, more resources allocated to $w_{1}$ yields lower smaller increase in $\varepsilon$, leading to $\tau_{w_{1}}>\tau_{w_{2}}$.
Situation 3.
In this third situation, the expected error $\varepsilon$ monotonically decreases in $\tau_{w}$ for one input, but is a non-monotonic function for the other.
Figure A3: Situation 3 expected retrieval error $\epsilon$ for two input words as a function of encoding precision $\tau_{w}$ and their surprisal.
Recall that whether $\varepsilon$ is monotonic depends on the inequality in Eq. A.28. That is, in order for $\varepsilon$ to be non-monotonic, Eq. A.28 must hold. Moreover, as discussed in Situation 2 above, if $h_{w_{1}}>h_{w_{2}}$, the turning point of monotonicity for the high surprisal input $w_{1}$ is to the left of the one for the low surprisal input $w_{2}$. As a result, in this Situation 3, it must be the case that it is the high surprisal input $w_{1}$ that monotonically decreases in $\tau_{w}$ whereas the low surprisal input $w_{2}$ first increases and then decreases, as illustrated in Figure A3.
Apparently, Situation 3 is basically equivalent to the Phase II and Phase III in Situation 2. Therefore, as proved above, if $h_{w_{1}}>h_{w_{2}}$, the optimal encoding should satisfy $\tau_{w_{1}}>\tau_{w_{2}}$.
Remark.
To sum up, if the surprisal of two input words $h_{w_{1}}>h_{w_{2}}$, given fixed amount of memory resources such that the encoding precisions for two inputs is constrained by $\tau_{w_{1}}+\tau_{w_{2}}=c$, the optimal encoding with strategic resource allocation should satisfy $\tau_{w_{1}}>\tau_{w_{2}}$ in order to achieve minimal total expected error $\varepsilon$. We outlined three possible situations of how $\varepsilon$ may change with increasing encoding precision $\tau_{w}$, and we proved that the strategic resource allocation should hold in all three situations. It is worth noting that, in most cases, $\varepsilon$ monotonically decreases with increasing encoding precision $\tau_{w}$, as in Situation 1. However, when the prior precision is too unreliable or when the input word is too close to the prior predicted word, there will be borderline cases where $\varepsilon$ first increases with increasing $\tau_{w}$ before it starts to decrease, as in Situation 2 and 3.
Appendix B Statistical models
B.1 Study 1
In Study 1, for each language, we ran regression models on dependency length $L$. As mentioned in the main article, the regression models were run separately for $L$ measured as intervening word counts and as intervening surprisal, as shown below in (1) and (2). For the analysis on full dataset, we ran linear mixed-effect model with random intercept per dependency type. For analysis on subject relations and object relations, we ran the standard linear regression without specifying random effects. Compare to the orthographic $L$, the analysis with information-theoretic $L$ includes baseline surprisal as an additional control variable.
  1. (1)
    Regression formulas for orthographic $L_{\mathrm{O}}$
    • Full dataset
      $L\sim$ 1 + Sentence Position + Antecedent Position + Sentence Length + Antecedent Frequency + Antecedent Surprisal + (1 $\mid$ Dependency Type)
    • Subject/object relations
      $L\sim$ 1 + Sentence Position + Antecedent Position + Sentence Length + Antecedent Frequency + Antecedent Surprisal
  1. (2)
    Regression formulas for information-theoretic $L_{\mathrm{I}}$
    • Full dataset
      $L\sim$ 1 + Sentence Position + Antecedent Position + Sentence Length + Baseline Surprisal + Antecedent Frequency + Antecedent Surprisal + (1 $\mid$ Dependency Type)
    • Subject/object relations
      $L\sim$ 1 + Sentence Position + Antecedent Position + Sentence Length + Baseline Surprisal + Antecedent Frequency + Antecedent Surprisal
B.2 Study 2a
In Study 2a, we ran linear mixed-effect models on log-transformed reading times for the critical region at the retrieval site (i.e., the right codependent for each syntactic dependency) and its spillover region.
  1. (3)
    Regression formulas for the critical region
    1. a.
      Fixed effects
      logRT $\sim$ 1 + sent.pos + antec.pos + sent.len + word.len + antec.freq + surp + surp.prev1 + surp.prev2 + freq + freq.prev1 + freq.prev2 + $L$ * antec.surp
    2. b.
      Random effects
      • Orthographic $L_{\mathrm{O}}$
        • Full dataset: (1 $\mid$ dep.type) + (1 $\mid$ part)
        • Subject relations: ($L$ + antec.surp $\mid$ part)
        • Object relations: ($L$ + antec.surp $\mid$ part)
      • Info-theoretic $L_{\mathrm{I}}$
        • Full dataset: (1 $\mid$ dep.type) + (1 $\mid$ part)
        • Subject relations: (antec.surp $\mid$ part)
        • Object relations: ($L$ + antec.surp $\mid$ part)
  1. (4)
    Regression formulas for the spillover region
    1. a.
      Fixed effects
      logRT $\sim$ 1 + sent.pos + antec.pos + sent.len + word.len + antec.freq + surp + surp.prev1 + surp.prev2 + freq + freq.prev1 + freq.prev2 + $L$ * antec.surp
    2. b.
      Random effects
      • Orthographic $L_{\mathrm{O}}$
        • Full dataset: (1 $\mid$ dep.type) + (1 $\mid$ part)
        • Subject relations: ($L$ * antec.surp $\mid$ part)
        • Object relations: (antec.surp $\mid$ part)
      • Info-theoretic $L_{\mathrm{I}}$
        • Full dataset: (1 $\mid$ dep.type) + (1 $\mid$ part)
        • Subject relations: (antec.surp $\mid$ part)
        • Object relations: (antec.surp $\mid$ part)
B.3 Study 2b
In Study 2b, we ran linear mixed-effect models on first-pass durations and total reading times at the retrieval site, as shown below in (5) and (6).
  1. (5)
    Regression formulas for first-pass durations (same maximal converging random structure for both measures of $L$)
    1. a.
      Fixed effects
      logRT $\sim$ 1 + sent.pos + antec.pos + sent.len + word.len + antec.freq + surp + surp.prev1 + surp.prev2 + freq + freq.prev1 + freq.prev2 + $L$ * antec.surp
    2. b.
      Random effects
      • Full dataset: ($L$ * antec.surp $\mid$ dep.type) + ($L$ * antec.surp $\mid$ part)
      • Subject relations: ($L$ * antec.surp $\mid$ part)
      • Object relations: ($L$ * antec.surp $\mid$ part)
  1. (6)
    Regression formulas for total reading times (same maximal converging random structure for both measures of $L$ in the analysis of full dataset and subject relations)
    1. a.
      Fixed effects
      logRT $\sim$ 1 + sent.pos + antec.pos + sent.len + word.len + antec.freq + surp + surp.prev1 + surp.prev2 + freq + freq.prev1 + freq.prev2 + $L$ * antec.surp
    2. b.
      Random effects
      • Full dataset: ($L$ * antec.surp $\mid$ dep.type) + ($L$ * antec.surp $\mid$ part)
      • Subject relations: ($L$ * antec.surp $\mid$ part)
      • Object relations:
        • Orthographic $L_{\mathrm{O}}$: ($L$ * antec.surp $\mid$ part)
        • Info-theoretic $L_{\mathrm{I}}$: ($L$ + antec.surp $\mid$ part)